A 4.0L flask of argon gas at 500 torr is attatched with a rubber seal to a 3.0L flask of nitrogen gas at 200 torr. What is the final pressure of this system assuming that this scenario takes place at room temperature, 25ºC.
1
Expert's answer
2017-03-26T04:53:06-0400
VNitrogen/VoArgon=7/4=1.75 Times increased the volume of argon
Pargon=500/1.75=285.71torr
Vnitrogen/VoNitrogen=7/3=2.3 Times increased the volume of nitrogen
Pnitrogen=200/2.3=85.71torr
P=285.71+85.71=371.47torr